OpenStudy (anonymous):

Hiram raises earthworms. In a square of compost 4 ft by 4 ft, he can have 1000 earthworms. How many earthworms can he have if his square of compost has a side length that is 8 times longer? A. 8000 B. 64,000 C. 256,000 D. 32,000

OpenStudy (anonymous):

i got A last time i did it, i mulitiplied 1000 by 8 , but it was incorrect :/

OpenStudy (anonymous):

(4 x 8)^2 / 4^2 = 1024/16 = 64 1000/1 = 64000/64 Answer: 64000

OpenStudy (anonymous):

thank you :)!

OpenStudy (anonymous):

Good luck toy you in all of your studies and thx for the recognition! And your're welcome!

OpenStudy (anonymous):

thanks again and no problem!